# SPIN-Bench: How Well Do LLMs Plan Strategically and Reason Socially?

### v3

@@ Abstract
Reasoning and strategic behavior in social interactions is a hallmark of intelligence. This form of reasoning is significantly more sophisticated than isolated planning or reasoning tasks in static settings (e.g., math problem solving). In this paper, we present Strategic Planning, Interaction, and Negotiation (SPIN-Bench), a new multi-domain evaluation designed to measure the intelligence of strategic planning and social reasoning. While many existing benchmarks focus on narrow planning or single-agent reasoning, SPIN-Bench combines classical PDDL tasks, competitive board games, cooperative card games, and multi-agent negotiation scenarios in one unified framework. The framework includes both a benchmark as well as an arena to simulate and evaluate the variety of social settings to test reasoning and strategic behavior of AI agents. We formulate the benchmark SPIN-Bench by systematically varying action spaces, state complexity, and the number of interacting agents to simulate a variety of social settings where success depends on not only methodical and step-wise decision making, but also conceptual inference of other (adversarial or cooperative) participants. Our experiments reveal that while contemporary LLMs handle basic fact retrieval and short-range planning reasonably well, they encounter significant performance bottlenecks in tasks requiring deep multi-hop reasoning over large state spaces and socially adept coordination under uncertainty. We envision SPIN-Bench as a catalyst for future research on robust multi-agent planning, social reasoning, and human–AI teaming.
Jianzhu Yao1∗, Kevin Wang2∗, Ryan Hsieh2, Haisu Zhou2, Tianqing Zou2, Zerui Cheng1,
Zhangyang Wang2, Pramod Viswanath1
1
Department of Electrical and Computer Engineering, Princeton University, New Jersey, US
2Department of Electrical and Computer Engineering, The University of Texas at Austin, Austin, US
{jy0246, pramodv}@princeton.edu{kevinwang.1839,atlaswang}@utexas.edu
* Equal contribution.
Website: https://spinbench.github.io
@@ 1 Introduction
However, few existing benchmarks provide a comprehensive, unified setting to rigorously test how LLMs balance these two intertwined requirements—strategic long-horizon planning and multi-agent social reasoning. Traditional AI planning benchmarks, such as those based on PDDL (Aeronautiques et al., 1998), usually emphasize step-wise problem solving for single agents under fully observable conditions. By contrast, conversation-focused benchmarks often evaluate an LLM’s linguistic coherence or role-play acumen (Li et al., 2023; Chen et al., 2023b) without imposing genuinely complex strategic demands.
To address these gaps, we introduce Strategic Planning, Interaction, and Negotiation (SPIN-Bench), a new evaluation framework specifically designed to capture both:
@@
    Strategic planning across single-agent formal tasks (PDDL) and highly interactive board/card games with large action and state spaces.
@@
Altogether, SPIN-Bench underscores both the real-world promise and current insufficiencies of LLMs in long-horizon, multi-agent tasks. We hope that our benchmark will motivate deeper research into strategic planning and social reasoning, paving the way for more capable human–AI partnerships and progress toward broader AI goals, including aspects of artificial general intelligence (AGI).
@@ 2 Related Work
2.1 Evaluation of LLM in Planning
LLMs have shown both promise and notable limitations in tasks demanding methodical, step-wise planning. Early work highlighted fundamental shortcomings in LLMs’ reasoning abilities, particularly on problem formulations grounded in PDDL (Hao et al., 2023; Valmeekam et al., 2023, 2024a; Wang et al., 2024). Yet, improvements in test-time scaling and prompt engineering have yielded models that can now achieve over 95% accuracy on discrete, place-based PDDL benchmarks (Valmeekam et al., 2024b; Wang et al., 2024). While these developments indicate an upward trajectory for LLM-based planners, current assessments tend to focus on relatively narrow or routine domains (e.g., classical planning formulations or daily schedule generation) (Zheng et al., 2024; Xie et al., 2024). Such specialized settings fail to expose many of the deeper challenges involved in long-horizon and multi-agent scenarios. To advance LLMs beyond these simpler planning tests, there is a need for more holistic benchmarks that illuminate how models handle large state and action spaces, manage partial or uncertain information, and adapt their strategies to non-trivial dynamics. These dimensions are especially vital if LLMs are to fulfill roles in complex decision-support systems and human–AI teaming.
LLMs for gaming Games offer an appealing testbed for LLMs because they often demand multi-step reasoning, strategic planning, and multi-agent interaction. Recent work on single-game evaluations has explored domains such as Minecraft (Gong et al., 2024), Avalon (Light et al., 2023), and Werewolf (Xu et al., 2023), but these narrow settings limit generalizability. Broader benchmarks incorporate multiple game types (Paglieri et al., 2024; Ruoss et al., 2024; Zhang et al., 2024; Duan et al., 2024b, a; Costarelli et al., 2024), and multi-agent frameworks emphasizing coordination or competition (Huang et al., 2024; Cross et al., 2024; Agashe et al., 2024), yet critical dimensions—like open-ended negotiation, evolving cooperation versus conflict, and richer social dynamics—remain underexplored. Other lines of research delve deeper into social intelligence in game-based contexts, investigating theory of mind (Cross et al., 2024; Hou et al., 2024), game-theoretic decision-making (Chen et al., 2023a; Akata et al., 2023; Abdelnabi et al., 2023), and dialogue-driven challenges (Qiao et al., 2023; Chen et al., 2024; Duan et al., 2024b; Abdulhai et al., 2023), but often within constrained strategies or limited forms of social dynamics. In contrast, our SPIN-Bench unifies strategic planning and social intelligence in a single testbed - combining formal PDDL analysis, multi-agent cooperation and competition, open-ended dialogues, and direct comparisons against optimal solvers and human baselines-to uncover richer insights into the interplay between step-wise action selection and conceptual inference in extended-horizon gaming environments.
@@ 3.1 Problem Definition: Three Settings for Action Sequence Generation
To capture the full spectrum of planning and social intelligence required in real-world scenarios, SPIN-Bench is organized around three progressively complex frameworks for automated action selection:

### v4

@@ Abstract
Reasoning and strategic behavior in social interactions is a hallmark of intelligence. This form of reasoning is significantly more sophisticated than isolated planning or reasoning tasks in static settings (e.g., math problem solving). In this paper, we present Strategic Planning, Interaction, and Negotiation (SPIN-Bench) [^1] Project page: https://spinbench.github.io/. Code: https://github.com/spinbench/spinbench, a new multi-domain evaluation designed to measure the intelligence of strategic planning and social reasoning. While many existing benchmarks focus on narrow planning or single-agent reasoning, SPIN-Bench combines planning domain definition language (PDDL) tasks, competitive board games, cooperative card games, and multi-agent negotiation scenarios in one unified framework. The framework includes both a benchmark as well as an arena to simulate and evaluate the variety of social settings to test reasoning and strategic behavior of AI agents. We formulate the benchmark SPIN-Bench by systematically varying action spaces, state complexity, and the number of interacting agents to simulate a variety of social settings where success depends on not only methodical and step-wise decision making, but also conceptual inference of other (adversarial or cooperative) participants. Our experiments reveal that while contemporary LLMs handle basic fact retrieval and short-range planning reasonably well, they encounter significant performance bottlenecks in tasks requiring deep multi-hop reasoning over large state spaces and socially adept coordination under uncertainty. We envision SPIN-Bench as a catalyst for future research on robust multi-agent planning, social reasoning, and human–AI teaming.
@@ 1 Introduction
However, few existing benchmarks provide a comprehensive, unified setting to rigorously test how LLMs balance these two intertwined requirements—strategic long-horizon planning and multi-agent social reasoning. To address these gaps, we introduce Strategic Planning, Interaction, and Negotiation (SPIN-Bench), a new evaluation framework specifically designed to capture both:
@@
    Strategic planning across single-agent formal tasks (PDDL) and highly interactive board games with large action and state spaces.
@@
Altogether, SPIN-Bench underscores current insufficiencies of LLMs in long-horizon, multi-agent tasks. We hope that our benchmark will motivate deeper research into strategic planning and social reasoning, paving the way for more capable human–AI partnerships and progress toward broader AI goals.
@@ 2 Related work
Evaluation of LLM in Planning LLMs show promise in methodical, step-wise planning but still face notable limitations. Early work revealed fundamental reasoning failures in PDDL-based tasks (Hao et al., 2023; Valmeekam et al., 2023; 2024a; Wang et al., 2024). Recent improvements in test-time scaling have enabled models to achieve over 95% accuracy on discrete PDDL benchmarks (Valmeekam et al., 2024b; Wang et al., 2024). However, these evaluations do not capture the deeper challenges of long-horizon planning. As outlined in Ghallab et al. (2004), planning involves understanding, reasoning, and execution; we further investigate at which stage each component begins to falter.
LLMs for gaming and social intelligence Games offer an appealing testbed for LLMs because they often demand multi-step reasoning, strategic planning, and multi-agent interaction. Recent work on single-game evaluations has explored domains such as Minecraft (Gong et al., 2024), Avalon (Light et al., 2023), and Werewolf (Xu et al., 2023), but these narrow settings limit generalizability. Broader benchmarks incorporate multiple game types (Paglieri et al., 2024; Ruoss et al., 2024; Zhang et al., 2024; Duan et al., 2024b; a; Costarelli et al., 2024), and multi-agent frameworks emphasizing coordination or competition (Huang et al., 2024; Cross et al., 2024; Agashe et al., 2024), yet critical dimensions—like open-ended negotiation, evolving cooperation versus conflict, and richer social dynamics—remain underexplored. Existing work builds a social AI data infrastructure (Li et al., 2024), and raises several core technical challenges to advance social intelligence (Mathur et al., 2024), while other lines of research delve deeper into social intelligence in game-based contexts, investigating theory of mind (Cross et al., 2024; Hou et al., 2024; Sclar et al., ), game-theoretic decision-making (Chen et al., 2023; Akata et al., 2023; Abdelnabi et al., 2023), and dialogue-driven challenges (Qiao et al., 2023; Chen et al., 2024; Duan et al., 2024b; Abdulhai et al., 2023; Zhou et al., 2024), but often within constrained strategies or limited forms of social dynamics. In contrast, our SPIN-Bench unifies strategic planning and social intelligence in a single testbed - combining formal PDDL analysis, multi-agent cooperation and competition, open-ended dialogues, and direct comparisons against optimal solvers and human baselines-to uncover richer insights into the interplay between step-wise action selection and conceptual inference in extended-horizon gaming environments.
@@ 3.1 Problem definition: three settings for action sequence generation
SPIN-Bench organizes automated action selection into three progressively complex frameworks — classical (single‑agent, deterministic) planning, multi‑agent (cooperative or competitive) games, and strategic games featuring dynamic alliances and negotiation. Each setting incrementally increases demands on an LLM’s planning and social reasoning abilities. Formal definitions, task formulations, and illustrative examples for all three settings appear in Appendix B.1. By exploring these three categories in tandem, SPIN-Bench systematically illuminates how LLMs grow from simpler single-agent planners into more complex negotiators and alliance builders, testing both mechanical planning skills and social inference in dynamic, multi-turn environments.
3.2 Game taxonomy and environments
PDDL.
Competitive games.
Cooperative game (Hanabi).
For multi-agent collaboration under imperfect information, SPIN-Bench leverages the cooperative card game Hanabi. Here, each player observes others’ cards but not their own, and must communicate implicitly to play cards in the correct order. This format demands not just straightforward planning but also trust-building, reasoning about other agents’ beliefs and intentions, and coordinated actions — essential for more advanced cooperative planning. Previous work (Bard et al., 2020) demonstrated that Hanabi requires advanced theory of mind reasoning, making it well-suited for evaluating social intelligence in collaborative settings. More details about how to play Hanabi are shown in Appendix B.4.
Strategic game (Diplomacy).
Finally, SPIN-Bench incorporates Diplomacy (Meta Fundamental AI Research Diplomacy Team et al.(2022)Meta Fundamental AI Research Diplomacy Team (FAIR), Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, et al., FAIR), a complex multi-player board game in which negotiation, dynamic alliances, and strategic betrayal are integral. Players exchange messages to coordinate or mislead, while simultaneously issuing movement orders to capture territories. Game introduction and tutorial are shown in Appendix B.5. Because negotiation is central, LLMs are tested on both long-horizon planning and social intelligence—managing alliances, inferring hidden intentions, and adapting their strategies in response to evolving alliances. This setting is especially suited for evaluating how chain-of-thought reasoning coexists with dynamic social interaction under partial cooperation and partial competition. It captures a broader spectrum of realistic multi-agent behavior that blends both cooperation and competition within the same session.
3.3 Benchmark construction
Classical planning data generation.
LLM game agent interface.
We implement a unified interface that supplies each model with the current state, relevant history, and a list of legal actions, alongside game-specific Chain-of-Thought (Wei et al., 2022) prompts. In Diplomacy, we integrate the Diplomacy game engine (Paquette et al., 2020) and include both public and private messages. Models may retry illegal moves up to ten times before incurring an automatic loss. Full details on engine integration, prompt templates and domain rules, appear in Appendix BE.
3.4 Evaluation metrics
Rule-based metrics.
For classical planning tasks, we measure both accuracy and N-Step Look Ahead (see Appendix B.2 for the formula). In competitive games, we compare LLM moves against solver-generated top-$k$ actions and track head-to-head outcomes against the solver. We also maintain a leaderboard of internal Elo ratings (do not directly reflect performance among humans). This helps us gauge the ”skill gap” among different LLMs and compare them against human baselines. For Hanabi, we use the final score to assess cooperative efficiency. Meanwhile, Diplomacy performance is evaluated through factual consistency, order-level success rates, and final game outcomes (see Appendix B.5 for details).
LLM-assisted negotiation metrics.
4 Experiments on SPIN-Bench
  1. 1.
  2. 2.
  3. 3.
In what follows, we first outline the overall experimental setup (Section 4.1). Next, we detail results on core planning skills (Section 4.2), examine the impact of action complexity and agent scaling (Section 4.3) and evaluate social intelligence and negotiation behaviors (Section 4.4). The reproducibility statement of all our evaluations is in Appendix A. Our trajectory viewer present game trajectories and examples in all benchmark settings.
4.1 Model and configurations
We evaluate a suite of popular LLMs that encompasses both closed-source commercial and open-source systems. The closed-source models include GPT-4o (Hurst et al., 2024), GPT-4o mini (OpenAI, 2024a), o1 (Jaech et al., 2024), o1-preview (OpenAI, 2024c), o1-mini (OpenAI, 2024b), o3-mini (OpenAI, 2025), o4-mini (OpenAI, 2025), GPT-4-turbo, GPT-3.5-turbo, Claude 3.5 Sonnet (Anthropic, 2024a), and Claude 3.5 Haiku (Anthropic, 2024b). Our open-source models include DeepSeek-R1 (Guo et al., 2025), Llama4 Maverick (MetaAI, 2025), Llama3-70b, Llama3.1-70b, Llama3.2-3b, Llama3.3-70b (MetaAI, 2024a; AI, 2024; MetaAI, 2024b; c), Qwen2.5-72b (Team, 2024), and Mistral-7b (Jiang et al., 2023). Since Diplomacy exhibits the largest branching factor in our benchmark, we first conducted a basic skill evaluation to identify which models should advance to the full Diplomacy task. Detailed results for the basic skill evaluation are provided in Appendix C.3.1. We excluded Diplomacy from the benchmark average score in Table 1 as few models play it well.
Classical Planning Competitive Games Collaborative: Hanabi Avg.
[…]
o1 58.59 16.09 70.0 0.0 0.0 83.1 45.9 16.4 14.8 14.8 14.2 49.8
o4-mini 46.79 11.52 80.0 0.0 0.0 81.1 50.5 12.8 11.0 12.6 13.2 45.6
o1-mini 13.20 1.95 50.0 0.0 0.0 87.0 36.5 6.8 7.4 11.4 10.2 33.0
o3-mini 51.25 13.04 20.0 0.0 0.0 74.2 52.8 8.8 7.6 8.8 8.0 33.1
GPT-4o 8.75 0.60 0.0 0.0 0.0 84.1 32.2 6.6 4.8 4.8 4.6 20.8
GPT-4-turbo 5.62 0.13 60.0 0.0 0.0 83.8 38.7 5.2 5.6 5.0 6.0 27.5
Claude 3.5 Sonnet 20.55 4.44 60.0 0.0 0.0 78.9 49.5 8.2 9.4 7.4 8.4 34.3
Claude 3.5 Haiku 4.22 0.30 50.0 0.0 0.0 69.6 35.9 2.4 4.0 2.8 2.8 20.8
DeepSeek R1 44.30 10.71 10.0 0.0 0.0 78.9 47.8 6.0 16.0 11.3 13.0 36.6
Llama4 Maverick 13.05 1.69 0.0 0.0 0.0 75.4 51.3 3.0 4.2 4.4 5.6 20.8
Llama-3.3-70b 5.78 0.32 0.0 0.0 0.0 79.5 25.4 2.4 0.8 0.6 1.4 13.1
4.2 Core planning competencies: factual retrieval,spatial reasoning and constraint following
We first examine why LLMs might falter on long-horizon planning by isolating two crucial facets: (i) factual retrieval and (ii) spatial reasoning. In addition, we perform an error analysis that highlights common failure modes.
Factual retrieval.
Figure 2: Heatmap displaying the F1 scores across evaluation categories in Diplomacy.
In Diplomacy, we design and categorize several factual queries into one-hop (e.g., ”Which locations does Russia control?”) vs. multi-hop (e.g., ”Which territories can France attack, and how many units of support are needed at least?”) to further check models’ factual retrieval in a highly strategic environment. As summarized in Figure 2, nearly all LLMs do well on basic location or adjacency checks but degrade by a large margin on ”Attackable” and ”Attack Analysis,” which demand deeper, multi-hop inference. Again, o1 and o1-preview lead but still exhibit significant drops compared to simpler tasks. We also examined the factual retrieval performance of models within the classical planning task in Appendix C.1. Our evaluation indicates that recent reasoning models can achieve accuracy exceeding 90%, while other models exhibit significant performance degradation as problem complexity increases.
Spatial reasoning.
To investigate whether LLMs’ planning deficits stem from weaker spatial understanding, we designed tasks based on PDDL format that require each model to track positions across sequences of relative movements. In these tasks, the LLM is given an initial spatial coordinate (e.g., (2,2)) along with a series of movements(up, down, left, right), and is tasked with computing the final coordinate. We evaluated all models on 90 scenes, each with a varying number of steps, to assess their spatial reasoning capabilities. Figure 3(c) plots the accuracy of each model against the length of the movement trajectory. Notably, o1-mini and GPT-4o exhibit declining performance as the number of steps increases, whereas o1 sustains perfect accuracy (100%) up to 29 steps.
Error analysis.
In classical planning, we categorized the errors into three distinct types: (1) Breaking Constraints (BC) includes execution failures, type mismatches, and structural issues, high BC reflecting difficulty in executing the reasoning; (2) Goal Not Satisfied (GS) refers to plans that satisfy all constraints yet fail to achieve the goal state; and (3) Others encompasses parsing errors, format inconsistencies, and miscellaneous validation issues. The results reveal a striking contrast: while Claude 3.5 Sonnet achieves a relatively low BC error (28.59%), it suffers from an exceptionally high GS error (44.77%). In contrast, DeepSeek R1—employing reinforcement learning–based training—drastically cuts the GS error to just 3.65%. This highlight that RL-based training LLM is particularly effective at reasoning, but failed in execution. The detailed quantitative breakdown is provided in Table 2.
[…]
o1 17.97 17.89 5.55
o1-mini 68.91 5.94 11.95
GPT-4o 81.56 3.59 6.09
Claude 3.5 Sonnet 28.59 44.77 6.09
Claude 3.5 Haiku 72.03 15.23 1.76
DeepSeek R1 48.28 3.65 1.87
Table 2: Error breakdown by model for the total of 1,280 problems.
In competitive board games, illegal‑move rates scale sharply with complexity: minimal in Tic‑Tac‑Toe, moderate in Connect Four, and highest in Chess (Appendix C.2), underscoring that rule compliance becomes a critical bottleneck as branching factors grow. In Diplomacy, o1 achieves the highest basic order success rate—20–30% above other models—but all models’ performance deteriorates on multi‑step and multi‑agent actions (e.g., support moves), revealing new failure modes rooted in deeper strategic and alliance reasoning (Table 3).
[…]
Figure 3: Relationship between accuracy and action space complexity. (a) Accuracy versus average number of legal actions for o1 model. (b) Accuracy versus grounded actions for o1 model. (c) Spatial reasoning accuracy across different models, where Claude refers to Claude 3.5 Sonnet.
4.3.1 Complex or expansive action spaces
PDDL domains.
We further examine the results of 10 domains in classical planning tasks by analyzing two key measures: the average number of legal actions available at each step and the total state–action space, defined as the complete set of possible state–action pairs given an initial state and action space. To evaluate their impact on model performance, we calculate the correlation between accuracy and these two metrics. As illustrated in Figure 3(a) and (b), accuracy demonstrates a stronger negative correlation with the total state–action space compared to the average number of legal actions. This implies that cognitively, the model ”carries” the burden of large future branching factors, even if only a few choices are valid at each step.
Competitive board games.
Across all three games, solver-level engines never lose, and LLMs fall short of optimal play (Table 1; full results in Appendix 8). In Tic‑Tac‑Toe, advanced models (e.g., o1, GPT‑4‑turbo, Claude 3.5 Sonnet) occasionally force a draw, but even here they underperform the perfect solver.
Model TTT C4 CH
Human 1415 1692 -
o1-preview 1263 1377 1395
o1-mini 1205 1083 1247
Claude 3.5 Sonnet 1138 942 1196
GPT-4o 1025 980 1255
Claude 3.5 Haiku 983 907 1041
Qwen2.5:72b 944 877 1150
Llama3.3 944 878 1146
Llama3.1:70b 936 829 1164
GPT-4o-mini 912 925 842
GPT-4-turbo 857 1007 1265
Llama3:70b 856 993 1155
Mistral:7b 855 800 243
Llama3.2:3b 836 821 421
GPT-3.5-turbo 829 890 480
Table 4: Internal Elo Ratings of LLMs (Initial Elo: 1000)
In Connect Four and Chess, every LLM lost $100\%$ of matches against solver-level engines. Top k action percentage analysis in Appendix(Figure 7) shows that while LLMs sometimes pick optimal moves in Connect Four, their accuracy drops drastically in Chess, underscoring how deeper tactics and branching expansions are beyond current LLMs’ capacity. Internal Elo ratings(Table 4) rank o1-preview and o1-mini highest, demonstrating strongest strategic planning capabilities, further indicating the advantages of test time scaling in strategic games. Additionally, Claude 3.5 Sonnet and GPT-4o follows, indicating a superior ability to handle complex strategies compared to their counterparts, yet all remain faw below solver or human-expert performance. These results highlight a significant gap in strategic reasoning and search depth, underscoring current LLM limitations in planning over large action spaces.
4.3.2 Scaling number of agents
In Hanabi, we vary the number of players from $2$ to $5$ and track final scores. Table 1 indicates that o1 remains the top performer, but its average score still declines from 16.4 (2-player) to 14.2 (5-player). This dip suggests that coordinating beliefs and actions among more agents still strains strong LLMs’ capacity to maintain consistent strategies over multiple incomplete information channels. However, models such as o1-mini and DeepSeek-R1 display erratic performance. Their high variance and relatively low average scores suggest insufficient strategic reasoning capabilities for effective Hanabi play. In Diplomacy, we vary the agent number from $2$ to $5$. Detailed results of more multi-agent settings are shown in Appendix C.3.2. As the agent count grows (beyond 2-3 test seats for LLMs), we observe decreasing order accuracy, fewer successful attacks, and minimal supply-center gains once multi-party negotiations and hidden alliances come into play. Ultimately, LLMs lose traction in highly interactive scenarios, underscoring how partial observability and shifting alliances further intensify the multi-agent complexity.
4.4 Social intelligence and negotiation performance
4.4.1 Cooperative scenarios vs. human baselines in Hanabi
Figure 4: Hanabi score distribution by player count (54,977 games)
We collected 54,97754{,}977 human-played Hanabi games from BoardGameGeek, spanning 2- to 5-player settings. Figure 4 plots the human score distribution, highlighting quartiles (Q1–Q4) around a typical range of 15–25 points. While some LLMs do show patterns of declining performance with more agents (cf. Table 1), none approach even the first quartile of human scores. This underscores the significant gap in cooperative planning under hidden-information constraints—despite Hanabi’s narrower branching factor relative to some competitive games. The results suggest a pervasive deficiency in social intelligence when LLMs must coordinate multiple incomplete information channels and track teammates’ evolving knowledge states.
4.4.2 Negotiation in Diplomacy
Overall impact on strategic play.
Negotiation phases in Diplomacy, which theoretically enable players to form or dissolve alliances, often produce counterintuitive effects on strong planners like o1, o1-preview. As shown in Table 3 and Table 10, o1’s all kinds of order numbers drop when negotiation is introduced, whereas other models sometimes stay the same or even improve slightly. And final game scores Supply Centers and Controlled Regions for o1, o1-preview also dropped by a large margin when enabling the negotiation, compared to other models. This result suggests that intense social interaction can disrupt planning coherence in otherwise capable LLMs, pointing to a tension between extended chain-of-thought reasoning and the cognitive overhead of real-time alliance-building, deception, or persuasion. In DeepSeek-R1, enabling negotiation features did not result in a significant performance decline. We attribute this to the model’s inherent tendency to propose one-sided solutions and provoke conflict during interactions (see the next section).
Negotiation message analysis.
We probe each model’s negotiation style in greater depth by prompting Claude 3.7 Sonnet to annotate 1,8641{,}864 messages across six-agent Diplomacy settings using six LLM-assisted metrics (see Appendix B.5.2 for details of the metrics; see Appendix C.3.3 for our human validation experiment of LLM annotation.). The results (see Table 3) enable the following key observations:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
5 Limitations.
A primary limitation of this work lies in the use of pre-defined prompt templates. While we conducted detailed ablation studies on prompt design effects in our most complex environment Diplomacy (Appendix E.3), the static nature of these templates may not fully elicit models’ latent social reasoning capabilities. Additionally, despite SPIN-Bench’s breadth across multiple game-theoretic environments, it cannot exhaustively capture the complete spectrum of real-world strategic and social interactions. Our analytical scope deliberately focuses on two core dimensions of social intelligence (Goleman, 2006): Theory of Mind, which corresponds to social awareness, and negotiation, which exemplifies social facility. This targeted approach, while enabling deep analysis of these critical components, necessarily excludes other fundamental aspects of social intelligence, including primal empathy, social attunement, persuasion dynamics, and self-presentation strategies. Consequently, the generalizability of our findings to broader social reasoning contexts remains bounded. Future work should explore adaptive prompt generation techniques that dynamically adjust to model capabilities, expand SPIN-Bench to encompass additional social intelligence facets, and investigate longitudinal multi-round interactions that better approximate the temporal dynamics of real-world social intelligence.
6 Conclusion
In this paper, we introduced SPIN-Bench, a comprehensive benchmark designed to assess strategic planning and social intelligence in Large Language Models (LLMs) across multiple game environments of escalating complexity. Our experiments encompassed formal PDDL-based planning, competitive board games, cooperative incomplete-information scenarios, and negotiation-intensive strategic settings. By systematically varying action and state-space size, interaction modalities, and the number of agents, SPIN-Bench exposes critical limitations in today’s LLMs. Despite recent improvements—particularly in short-range planning and factual recall—our findings reveal that most LLMs struggle with long-horizon tasks involving large branching factors and intricate multi-agent coordination. We also observe that negotiation and social interplay often degrade an LLM’s chain-of-thought coherence, suggesting a tension between pure strategic reasoning and the cognitive overhead of dynamic alliances and potential deception. Overall, our work highlights pressing gaps in both deep, multi-hop reasoning and robust social interaction, pointing toward the need for novel architectural innovations, integrated planning modules, and more advanced training methodologies. We hope SPIN-Bench will serve as a foundation and catalyst for continued progress in developing LLMs that are more strategically competent, socially adept.
Appendix A Reproducibility statement
Compute.
Appendix B Experiment settings
B.1 Problem definition: three settings for action sequence generation
Figure 5: Sample visualization from our PDDL visualizer.
To aid rule understanding, we’ve included animations for PDDL and interactive games. Try them here:: https://spinbench.github.io/tools/pddl/tra.html